# Bridging the Social & Technical Divide in Augmentative and Alternative Communication (AAC) Applications for Autistic Adults

### v1

@@ 5.7.3. Limitations of showing text
Despite showing text being commonly used and even used as a back-up for TTS-based output, there are some severe limitations to it that auditory communication does not have. Screens are often too small to show text (1), which makes it particularly difficult to share with people at a distance (1) or sharing something with a group of people (2). Because of this, especially with phones, users have to hand their device to another person (2), which can be a security risk. Reading from a smart device can have extra complications since there can frequently be a glare on the screen, especially when outside (1).
@@ 5.7.4. Device restrictions
There are times when our participants (6) struggled with the physical devices themselves. This is becomes obviously when an AAC application is put onto a regular phone (vs a tablet). A phone’s screen is typically too small to see the symbols (2), but even using text-based AAC requires the user to scroll more to see text output (1), which is difficult when showing text to someone. One participant (1) wears headphones a lot, which are connected to their phone, and so it’s not easy to switch the audio to let others hear the synthesized voice from an AAC app. Another participant (1) finds the screen so limiting that they wish there was a computer program for their app so that they could edit boards on their PC before putting them on their phone.
One of our participants (1) has a particularly interesting device called the UbiDuo—shown in Figure 3, which has a standard keyboard and is not connected to the internet. However, it is incredibly bulky and single-purpose, unlike a tablet or phone.
I mentioned it was bulky, which created a lot of barriers. Number one, the weight, and then not everything fit in my backpack, I couldn’t have it on my desk and do my work. So, I’d have to get it up and down to use it.
Even tablets aren’t perfect as older devices can have unpredictable quirks (1), such as crashing in extreme temperatures (1).
@@ 5.7.5. Unreliability
One point that was repeatedly brought up was the wish for continued, reliable use for AAC (6). It’s so important for one participant (1) that: “I always have and always will carry pen and paper just because it’s the most reliable”. It’s especially important when even commonly-used applications have bugs. Several participants (5) ran into bugs such as sound glitches (1), external notifications breaking off TTS (1), or erasing what the user was typing (1). Although they didn’t mention bugs explicitly, one participant’s (1) device couldn’t say ”TTY” during the interview and they had to type it in the chat. It can be disheartening, as well, when bugs are reported but the request to fix them is dismissed.
@@
5.7.6. Language choices
Most AAC is English-specific (Amery et al., 2022). People who speak multiple languages or are learning another language are stuck having to find another AAC application to speak in that language (Swedish; 1), attempting to make their own (Finnish; 1), or never talking in that language (Irish; 1).
In particular I’ve put about two years into learning Finnish but still can’t get any AAC apps in Finnish because [it’s] not like there’s many people who want to deal with any non Indo-European languages, but I know so many people who’ve run into this—including with Indo-European languages that are not popular languages.
5.7.7. External support
Customer support can be key to whether or not a user will keep using an application. Participants who experienced good customer support (2) feel like they can rely on the company to continually improve the app. On the other hand, participants who experienced bad customer support (3) will switch applications, if possible, or suffer in silence since their concerns are not being actively addressed by the developers.
Beyond the company themself, users might be able to find support through tutorials online (1), but less commonly-used applications have fewer users and therefore less of a community to draw from (1).
I wish there was […] an export feature where […] there could be a forum, so that I could share my board, and you could share your board, and everybody could.
5.7.8. Restrictive features
Restrictive features (1) or attributes of an app that make it harder to customize (2) can be significant hindrances. Occasionally, these can be okay if they can be turned off or unused without hurting the interaction with the app (1). A couple examples of restrictive features that people reported include having people unable to call someone back so that they can use Relay (1) and having to use the TTS for pre-programmed phrases because they don’t show up in the text box when selected (1).
5.8. Social Concerns
Many social issues arose even when prompting participants about how AAC technology can be improved. All of these social concerns can be boiled down to a lack of respect from conversation partners and other people in the user’s environment. Most of the time this lack of respect stems from ignorance. The majority of our society is not used to seeing people use AAC and therefore do not know how they are supposed to react. Many times if people had just acted normally it would have been better than doing what they believe is “accommodating”.
5.8.1. Communication not being taken seriously
Many participants (8) spoke about situations where they were not taken seriously by conversation partners. A few participants (4) have run into others treating them as if using AAC means they’re less capable of making decisions or being competent. In particular, there is a perceived stigma and infantilization around using symbol-based AAC (2)—“people seem to think symbols means you’re less capable or something”.
Meanwhile, others were concerned they wouldn’t be taken seriously had they used their mouth-words (1) or had switched back and forth between using AAC (1). For example, one participant (1) had used AAC in the past to make sure they made a good initial impression on their classmates and instructor: “I wanted them to take me seriously, and I knew that my voice would not come out clean and fluent and that would be held against me”.
5.8.2. AAC use not normalized
All of our participants (12) spoke about how AAC use is not normalized in society and described social situations where they encountered people not accepting or expecting AAC as communication. Participants will sometimes have to explain or have others explain for them that they’re switching to AAC (2) or will avoid using AAC because they don’t want to deal with explaining it to every person they run into (1). Occasionally people don’t understand why AAC is being used (2) or at least don’t expect it (2). Many people do not know how to communicate with AAC users and will assume someone who cannot speak orally is deaf or hard-of-hearing (4). Conversation partners will also, subconsciously or consciously, gesture more often (1) or use fewer mouthwords themselves (1) when speaking to an AAC user.
In the worst cases, conversation partners will not be respectful (4) or will outright not accept its use (3). Participants who occasionally use mouth-words will get weird responses when they need to switch to AAC at some point (3). Therefore, participants think there needs to be more education around how to respect AAC users (4) — “there are some cases where I need to train people to respect AAC, and I’m not sure about how to manage that yet”.
5.8.3. Prevented from using AAC
Due to the lack of normalization, several instances came up where participants were prevented from speaking because their AAC options were taken away (6), such as their whole device being taken (4). Two participants had their devices confiscated at hospitals because the device either had a camera in it (1) or because it was a “no phone space” (1).
One participant (1) had gotten around this in a courtroom by using low-tech letterboards, but even low tech options are not immune to being removed as an option. Another participant (1) spoke about how breaking their right arm in 7th grade—where writing was their primary mode of communication—meant an entire 6 months where they could barely communicate.
Occasionally devices are taken away so that others could read what they had written (1) or to even play around with the application (1). One participan t(1), who uses the bulky UbiDuo, recounted a harrowing ordeal at a basketball game they went to:
They search any bags you bring, and you have to walk a long ways to get there, so it didn’t make sense to try and bring a device outside of a phone. […] A UbiDuo would be too big and then they might freak out cuz it’s technology. So I went to this event, and it ended up being really loud. They had a loud halftime show, and I lost speech and a lot of motor skills. I needed a different way to communicate, and I couldn’t because I didn’t have the right tools. My hands would not cooperate for typing on a phone. My body was jerking around enough that they couldn’t read my signs.
5.8.4. Impatient conversation partners
Many participants struggle with impatient conversation partners (7) or have trouble getting people to listen/pay attention (4). AAC users will often have others talking over them (2), especially in groups (4). Occasionally people will try to guess what the AAC user is about to say in order to speed things up (1), but this will often make the conversation slower.
I remember going to meetings of the disability group I started in my school, and they just all kept talking so quickly and changing topics that it was frustrating to try to participate with AAC. And I felt like I either had to give in and use mouth words, or I wouldn’t be able to say what I was thinking”.
@@ 5.9. Control of Communication
Wanting to be autonomous and in control of their own communication, autistic adults have concerns about people and companies potentially having access to everything they say. Many AAC applications are made with SLP intervention in mind, as children are learning to speak for the first time with their AAC. However, adults finding AAC later in life do not need these supports and the conventional logging practices can seem invasive at best. We begin this section with a discussion about being in Control of features, where we will also dive into the privacy ratings we collected at the end of the interview and cross-examine them with our qualitative analysis. Unwanted outsider access and Company trust also arose as themes in the context of privacy.
@@ 5.9.1. Control of features
In addition to the semi-structured questions, we listed features that loosely aligned with modern natural language processing techniques and asked participants whether they would use them, with the framing of privacy (never, only in the current conversation, across conversations, or always okay). The features were: automatic speech recognition, logging conversations verbatim, logging the topic of the conversation (i.e., topic modeling), saving personalized information automatically, saving personalized information explicitly, using a history of sentences said in the past (i.e., language modeling), and using a history of words said in the past (i.e., bag of words). The full table that the participants saw can be found in Appendix B.3.
The results of the responses, in aggregate, are shown in Figure 4. Overall, deliberate personalization was the most accepted, which makes sense since a lot of our participants take the time to personalize their AAC with words that are unique to them. This type of information seems to be mostly okay for applications to store. Although the question of whether it’s stored locally or in the cloud might change their use.
Figure 4. Level of comfort for use of and data storage from 7 potential features. Participants rated each feature as either something they would never use (red), something they would only use as long as the data persisted only in their current conversation (orange), something they would use with the data persisting across conversations — with the assumption that it would be deleted over time (yellow), or something they would be okay using all the time (green). The specific descriptions for each feature can be found in Appendix B.3.
[…]
Number of people who would use each feature. Automatic personalization: 6 never allow, 2 current conversation, 3 across conversations, 1 okay Speech recognition: 5 never allow, 4 current conversation, 1 across conversations, 2 okay Log verbatim: 2 never allow, 4 current conversation, 6 across conversations, 0 okay Log topics: 2 never allow, 1 current conversation, 8 across conversations, 1 okay Sentence prediction: 1 never allow, 2 current conversation, 5 across conversations, 4 okay Word prediction: 1 never allow, 0 current conversation, 5 across conversations, 6 okay Deliberate personalization: 1 never allow, 0 current conversation, 2 across conversations, 9 okay
While talking about privacy (starting at Question 6), some participants (4) mentioned how they were okay using features as long as they’re explicitly used and can be turned off easily. People don’t want to be forced into certain features (1) and they want the ability to have control over details of how a feature is used (1).
@@
Further, automatic personalization was the least liked in our quantitative analysis (Figure 4). Half of our participants said that this was never a good feature, more than any other feature.
Some participants would even occasionally allow invasive data collecting such as the logging of conversations as long as they had control over what it was logging. For example, if there was a way to delete items (1) or to make sure certain conversations aren’t logged (1). Others would never allow logging that couldn’t be turned off (1). Still others were fine with logging as long as it didn’t analyze their data (1). However, none of our participants were comfortable having logging verbatim always on (Figure 4).
I turned off every single prediction that ClaroCom, including its built-in support for learning automatically. I don’t even like that it has those knobs. I’d rather it just plain didn’t do any of that. But unfortunately, needs must.
Speech recognition also had 5 ”never” ratings, perhaps because of the thought of the app “listening in” on other conversations or on conversation partners who did not consent to be transcribed.
@@ 5.9.2. Unwanted outsider access
Although some participants find company data collecting useful, participants do not want this collection to happen automatically. Most participants (11) were concerned about outsiders either getting a hold of their data without their consent. Most of the time this concern is centered around companies logging data. While data security was certainly a concern, participants expressed explicit concerns about logs of conversations ending up in the wrong hands—such as authorities (1), parents (1), or “others” such as SLPs (1). The concern being that others using these logs can use them to abuse the AAC user (1).
The concerns were raised about sensitive data being saved. One participant (1) wanted to make sure that they had legal access to the things that they say. Another participant (1) was less concerned about their own words being seen and more concerned about the information they share with other people. At their job, they had to take extra measures to make sure that their AAC was HIPAA-compliant:
I do have to think a little about data security specifically for my job because I’m bound by HIPAA and some stuff I say is confidential, but the HIPAA compliance person at my job, talked through that with me to make sure that things were in compliance.
People (2) were also concerned about tracking data that they felt was unnecessary for an AAC app to track, such as location. Other participants showed concern about people reading what they’re saying before they say it (2), other parties reading text intended for conversation partners (1), or conversation partners going through their phone when they need to show them the screen or share their device with them (2):
@@
Only one participant didn’t express immediate concern about outsider access. They were the only participant who has been using various types of AAC since childhood. They have used a variety of AAC technologies their whole life, most of which were created pre-internet and only had data stored on the device. Now, their AAC use involves writing on paper and using services such as Relay that have professionals reading their text aloud.
Well, they’re held a really strict confidentiality standards, so I never have concerns about that. […] [they] can lose [their] license […] And if the sensitive information I would be sharing on Relay would be medical stuff, it’s a stranger.
@@ 5.9.3. Company trust
Some participants (4) trust the company of their AAC applications well enough. One participant (1) can tell that the company of the app they use (AssistiveWare) appreciates feedback, and this helps them know that they care about their user base. However, even well-intended companies might slip up sometimes. One participant (1) told a cautionary tale of a company misusing data. That is, they thought they could use anonymized data without permission (1).
Many tech companies rely on data to train machine learning models, but even without training models, companies need to know how people are using their applications and how they can be improved. So what type of data are AAC applications saving about their users? Nine participants (9) mentioned some uncertainty about the data policy of their AAC application, either in general or in the specifics of particular features of the app. So it makes sense that people (2) would appreciate clear data policies. One participant (1) teaches people how to clear the cache for their AAC apps so that they can have some control over what their phone is storing.
6. Guidelines for Improving AAC Use
In this section, we discuss a set of recommendations based on our analysis for how AAC features can be improved from a research perspective, how AAC should be designed and implemented for autistic adults, and what policy makers might be able to do to improve access and normalize AAC use.
6.1. Guidelines for Natural Language Processing (NLP) researchers
As large language models (LLMs) like ChatGPT become more prevalent in AAC technologies, we need to be able to integrate these new methods while also keeping the user centered. In terms of what autistic adults need from the NLP research community, we have identified three major thrusts.
6.1.1. Faster communication
Although there has been a lot of work in word prediction (see Section 2), there is still a long way to go to integrate them smoothly into AAC applications so that there is no additional overhead. Even though LLMs can produce much more coherent text, they take a lot of compute power and time to run. Therefore, despite the potential for using LLMs in AAC, they are not a viable option when it comes to speed and the ability to run on a phone. If LLMs are going to be used, they must not be during a live conversation. Smaller models will have to be used for on-the-fly word prediction.
6.1.2. Transparent data collection
Machine learning models need data to train on, but nobody wants every single conversation that they’ve ever had to be recorded and put into a model for everyone to use. Therefore we need transparent data collect where we tell users exactly what will be recorded if they opt-in to data collection in the first place. It should be explained in as much detail as possible (and with no unexplained technical jargon) what the data could potentially be used for. Furthermore, there should be an obvious way for the user to temporarily not collect data and/or remove certain sentences from what was collected. Certain sensitive topics might end up with sparse data, but perhaps users can put up with less predictive for those specific conversations — as long as it is explained to them.
6.1.3. Voice customization
AAC devices should provide inclusive and diverse text-to-speech voices so that users can find a voice that matches their identity. They should have the ability to use a voice that is unique to them and can be used across multiple devices. The voices should be easy to understand, able to be played at a loud volume, and a realistic as possible.
6.2. Guidelines for AAC application makers
Here we will provide suggestions for anyone who wants to make AAC applications, regardless of whether you’re a company, a researcher, or an individual developer who wants to make their own application.
6.2.1. Transparency of features and data
First and foremost, it is essential for applications to be transparent with what data they save, what data they store locally on the user’s device, what data they store on a separate server, and if/how the data is being processed. All of this must be described in plain language (no tech/legal jargon) so that users can understand what to expect from the application and to build trust in the company who made the app. No data should be collected without the explicit permission of the user. Beyond this, any feature that is not considered part of the core functionality of the application must be “opt-in” rather than “opt-out”. That is, users must know that the feature is there so that they can decide whether or not to turn it on.
6.2.2. Modular design
Since all of our participants valued high levels of customization and a variety of features to fit their needs, we recommend that application makers adopt a modular design. Although the implementation is non-trivial, modular AAC applications would enable users to customize the app to their liking without sacrificing the needs of other users. This can be done in part by allowing for third-party developers to create plug-ins. These plug-ins would have to follow the transparency guidelines outlined in Section 6.2.1. This will encourage users to augment their AAC and share their modifications with the community. It will create standards across users so that even people with very different apps can share components they find useful without a large overhead.
6.2.3. Adult-first design
Many existing AAC applications are created with children in mind, and this leaves autistic adults with having to force applications to fit their needs. We suggest that there should be some AAC applications that are designed for adults first, since they are just as important of a group of users as children are. Furthermore, children will eventually grow up and will most likely want to shift to AAC that reflects their development. Design decisions like this might include less recording of button presses and less reporting functionality.
6.2.4. Variety
We would like to encourage more companies to create AAC applications. By having a variety of applications—in conjunction with making AAC use more commonplace, we can cover a larger suite of needs and drive down the price of individual app purchases. Once prices are cheaper, more people might be willing to pay for AAC.
6.3. Guidelines for policy makers
6.3.1. Greater support of use
We can support adult AAC users by providing more access to SLPs and other professionals who can setup AAC. These professionals should be educated on autistic adult perspectives so that they are informed before transitioning from child to adult patients or from allistic (not autistic) to autistic patients. There should also be a greater effort to make dedicated AAC devices free or affordable.
6.3.2. Broadcasting of existing accessibility programs
Participants spoke about situations where they weren’t aware of a certain feature/device being available to them. One participant highlighted that numerous users are unaware that smartphones often come with built-in features like Text Relay. Moreover, many AAC users are not informed about state government funds that they could access. It is essential to provide resources to educate AAC users or potential AAC users about the range of options already available to them and have a compilation of resources available depending on what state they live in.

### v3

@@ 5.7.4. Limitations of showing text
Despite showing text being commonly used and even used as a back-up for TTS-based output, there are some severe limitations to it that auditory communication does not have. Screens are often too small to show text (1) or can have a glare (1), which makes it particularly difficult to share with people at a distance (1) or with a group of people (2). Because of this, especially with phones, users have to hand their device to another person (2), which can be a security risk.
@@ 5.7.5. Device restrictions
There are times when our participants (6) struggled with the physical devices themselves. Smartphone screens are often too small to work for traditional AAC applications (3). One of our participants (1) has a particularly interesting device called the UbiDuo—shown in Figure 3, which has a standard keyboard and is not connected to the internet. However, it is incredibly bulky and single-purpose. Even tablets are not perfect as older devices can have unpredictable quirks (2), such as crashing in extreme temperatures.
@@ 5.7.6. Unreliability
One point that was repeatedly brought up was the wish for continued, reliable use for AAC (6). It is so important for one participant that they “always have and always will carry pen and paper just because it’s the most reliable”. It is especially crucial when even commonly-used applications have bugs. Several participants (6) have run into bugs while using their AAC. It can be disheartening, as well, when bugs are reported but the request to fix them is dismissed:
@@ 5.8. Control of Communication
How companies log and use data has long been a concern within AAC research (Pennington et al., 2007; Reddington and Coles-Kemp, 2011). Wanting to be autonomous and in control of their own communication, autistic adults have concerns about people and companies potentially having access to everything they say. Many AAC applications are made with SLP intervention in mind, as children are learning to speak for the first time using AAC. Adults finding AAC later in life might not need or have access to these supports, and the conventional logging practices can seem invasive at best.
At the end of each interview, we listed features that loosely aligned with modern natural language processing techniques and asked participants whether they would use them, with the framing of privacy (never, only in the current conversation, across conversations, or always okay). The features were: automatic speech recognition, logging conversations verbatim, logging the topic of the conversation (i.e., topic modeling), saving personalized information automatically, saving personalized information explicitly, using a history of sentences said in the past (i.e., language modeling), and using a history of words said in the past (i.e., bag of words). The table that was shown to participants can be found in Appendix B.3, followed by the results of the responses in aggregate (Figure 4).
@@ 5.8.1. Control of features
While talking about privacy (starting at Question 6), some participants (4) mentioned how they were okay using features as long as they are explicitly used and can be turned off easily. People do not want to be forced into certain features (1) and they want the ability to have control over details of how a feature is used (1).
@@
Further, half of our participants (6) said that automatic personalization was never a good feature, more than any other feature in our quantitative analysis (Figure 4).
Some participants (4) would occasionally allow the logging of conversations as long as they had control over what, how, and when it was logging — for example, if there was a way to delete items (1) or if the data was not analyzed in any way (1). However, none of our participants were comfortable having logging verbatim always on (Figure 4).
I turned off every single prediction that ClaroCom [has], including its built-in support for learning automatically. I don’t even like that it has those knobs. I’d rather it just plain didn’t do any of that.
Speech recognition also had 5 “never” ratings, with two participants (who both said “never”) saying that they would be interested in that only if it was outside of an AAC application, and a third participant (“current”) would use it “only if deliberately activated”.
@@ 5.8.2. Unwanted outsider access
Although some participants find company data collecting useful, participants do not want this collection to happen automatically. Most participants (11) were concerned about outsiders collecting their data without their consent. Mostly this concern was centered around companies logging data. While data security was certainly a concern, participants expressed explicit concerns about logs of conversations ending up in the wrong hands—such as people who hold control over the user (3).
People (2) were also concerned about tracking data that they felt was unnecessary for an AAC app to track, such as location. Other participants showed concern about people reading AAC text without their permission (3) or conversation partners looking through their phone (2):
@@
Only one participant (1) didn’t express immediate concern about outsider access. They were the only participant who has been using various types of AAC since childhood, many of which were created pre-internet and only had data stored on the device. Now, their AAC use involves writing on paper and using services like Relay which have professionals reading their texts aloud:
[Relay professionals are] held to really strict confidentiality standards, so I never have concerns about that. […] [they] can lose [their] license […] And if the sensitive information I would be sharing on Relay would be medical stuff, it’s a stranger.
@@ 5.8.3. Company trust
Some participants (4) trust the company of their AAC applications well enough to handle their data. Nonetheless, nine participants (9) mentioned some uncertainty about the data policy of their AAC application, either in general or in the specifics of particular features of the app, and a couple participants (2) desire clear data policies.
6. Research Directions
In this section, we highlight some of the technical aspects that our participants brought up which may improve AAC but have not been, to our knowledge, mentioned much or at all in previous literature. Of course, these guidelines should be treated as such: guidelines. Any AAC applications that are being made should be prototyped and tested with their target user base. Furthermore, while these guidelines were discovered from autistic adults, we believe many of these would aid other populations of AAC users as well.
6.1. Conversation Planning
Some of our participants considered the slower speed a feature of AAC, with participants using symbol-based AAC for calmer, more purposeful conversation (§5.1.2). They use AAC to collect their thoughts and remember what they said previously in the conversation (§5.5.1), or to queue up things to say while they wait for a pause in the conversation (§5.4.2). We recommend leaning into the unique capabilities of AAC by researching ways to view conversation history, pull up previous conversation topics (summarized conversation history), or plan conversation topics. Furthermore, how can AAC aid with conversation planning without logging conversation history (§5.8.2)?
6.2. Keyboard Input
Keyboard alternatives (e.g., (Black et al., 2019)), alternative keyboard mechanics (e.g., Wiegand and Patel (2014)’s Swype-like interface), and other novel modalities for AAC have been extensively researched (Ascari et al., 2021; Curtis et al., 2022, 2024). There is, however, surprisingly little work on regular keyboard use with AAC applications (Curtis et al., 2022).
All of our participants (§5.1.1) liked some form of keyboard functionality. Integrating a physical or on-screen keyboard made it so that users could leverage their typing speed or prevent the need to hunt for buttons. Proloquo includes an on-screen keyboard to allow the user to switch to typing in words, but only one word can be entered at a time. Some of our participants who used CoughDrop integrated a keyboard into their home board (§5.1.1, Figure 2) but future work should look into how AAC applications can intentionally integrate keyboards. An initial step would be to include Bluetooth capabilities within AAC apps so that users can connect a physical keyboard, but future work should investigate how to balance speed and convenience.
6.3. Opt-in Functionality
Any feature that is not considered part of the core functionality of the application should be “opt-in” rather than “opt-out” (§5.8.1). That is, users must know that the feature is there so that they can decide whether or not to turn it on. We need transparent data collection where users are told exactly what will be recorded only if they opt-in to data collection in the first place. Potential research directions here include (1) how to give the user control over when to collect data and/or remove certain sentences from what was collected; or how to best present and describe features so that users know they are available and know what they would do. For example, visualization techniques might be useful for showing a user how their data is collected (Hamidi et al., 2018), but how might this be extended to show what was collected or what a feature does?
6.4. Presentation Support
A few participants mentioned the use case of giving presentations using their AAC (§5.4.3). Due to the short, conversational form that AAC is usually designed for, users run into issues when using AAC for longer forms such as presentations. Research directions here might include:
  1. (1)
    Drafting paragraphs of text instead of single sentences
  2. (2)
    Making tablets ”project” TTS loudly enough or easily connect AAC applications to a speaker system
  3. (3)
    Coordinating AAC applications with slideshow software
  4. (4)
    TTS that is appropriate for emotive presentations (i.e., a presentation mode for TTS)
6.5. Same-app Switching
Beyond customizing an AAC application for a specific user, it should enable a user with dynamic disabilities to switch between features, depending on their current needs, without changing applications. We identify three examples that we believe would be worthwhile research directions for same-app switching. (1) Easily switching between direct and indirect input5.1.6), (2) simplifying the app’s colors or board when the user is overwhelmed5.3.3), and (3) combining symbol-based and text-based AAC within the same app (§5.1.2,5.3.1). The application should integrate these in an intentional way that keeps the user in control (§5.8.1) and does not impede motor plans (§5.1.1).
Separate from Keyboard Input, # looks at the ability to switch between symbol- and text-based AAC depending on the user’s needs. We have not found any evidence of research or products that attempt to switch between the two separate modalities. Integrating them into a single app could save the user time since they would not need to switch applications, and it would give them access to the same saved phrases across modalities.
6.6. Trans* Voices
While wanting personalized voices is not unique to autistic adults, there is a large overlap between trans* and autistic populations (Warrier et al., 2020), making gender-diverse speech synthesis valuable to the autistic community. Our participants want inclusive and diverse text-to-speech (TTS) voices so that they can find a voice that matches their identity, not necessarily their body’s voice (§5.2.2). Research should continue to look at creating voices not based on voice cloning and understanding how trans-feminine, trans-masculine, nonbinary, and genderqueer people sound and/or want to sound with TTS. There is some work on creating gender-expansive voices for TTS (Hope and Lilley, 2020; Székely et al., 2023; Hope, 2024; Szekely and Hope, 2024). In addition to improving these voices themselves, future research needs to look at how people feel when using these voices for their daily AAC and how to create TTS voices that can be customized by trans users.